OpenStudy (anonymous):

Find a formula for a linear function that satisfies the given conditions: an initial value of 3 and increases by 1/4 when x increases by 1

OpenStudy (anonymous):

What is the formula for a linear equation?

OpenStudy (anonymous):

Y=mx+b?

OpenStudy (anonymous):

Exactly. The initial value of the formula is 3. How does that fit into your equation?

OpenStudy (anonymous):

3=1/4x+1??

OpenStudy (anonymous):

Almost, but not quite. What is the value of x, when we have the value 3? (BTW, your final answer will be in the form y = mx + b, where m and b are replaced by numbers).

OpenStudy (anonymous):

Y=1/4(3)+1?

OpenStudy (anonymous):

Not quite. You can think of the initial value as the y-intercept of the equation.